OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -6, an = 4 • an-1.

OpenStudy (anonymous):

\(\Huge a_{\color{red}{n}} =4*a_{\color{red}{n-1}}\) if n =1, you have a1, right? and a1=-6 if n =2, you have a2, right? and \(\Huge a_{\color{red}{2}} =4*a_{\color{red}{2-1}}\) = what??

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

why 4? \(a_{2-1}=a_1=-6\) so, \(a_2=4*(-6)=-24\) you can find other terms by this way. I am done in helping. I can check,

OpenStudy (anonymous):

-6, -24, -96, -384, -1536, -6144 ? @OOOPS

OpenStudy (anonymous):

sounds like a geometric sequence ♥

OpenStudy (anonymous):

Haha thanks

OpenStudy (anonymous):

Yes, you are correct